OpenStudy (anonymous):

your brother earned 6 less than you did working. if you both earned a total of 60 dollars how much did each of you earn

OpenStudy (anonymous):

for this problem i would just guess. maybe 24 and 30?

OpenStudy (anonymous):

no they don't add to 60

OpenStudy (anonymous):

how bout 27 and 33?

OpenStudy (anonymous):

if you want an equation you could write \[n+n+6=60\] \[2n+6=60\] \[2n=54\] \[n=27\] and \[n+6=33\] but it is too much work

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw